The sum of three consecutive even numbers is 30. What is the greatest of the three even numbers?
Answer by checkley77(12844)

LET X, X+2 & X+4 BE THE THREE CONSECUTIVE EVEN NUMBERS.
X+X+2+X+4=30
3X+6=30
3X=30-6
3X=24
X=24/3
X=8 ANS. FOR THE SMALLEST NUMBER
8+2=10 FOR THE MIDDLE NUMBER
8+4=12 FOR THE LARGEST NUMBER.
PROOF:
8+10+12=30
30=30